OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent?

OpenStudy (anonymous):

The value of (x - 4) when x = 3 The value of x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

zepdrix (zepdrix):

The `function` is `(x-4)`. f(3) represents the value of the `function` when `x = 3`.

OpenStudy (anonymous):

so the first choice? :L

zepdrix (zepdrix):

Hopefully that helps you to make the connection to the correct answer D: I can make it a little bit more explicit if needed. \[\Large f(\color{#DD4747}{x})\quad=\quad (\color{#DD4747}{x}-4)\]\[\Large f(\color{#DD4747}{3})\quad=\quad (\color{#DD4747}{3}-4)\] So here we can see that f(3) represents the solution to (x-4) when we plug x=3 in.

zepdrix (zepdrix):

Yesss good good good :D

OpenStudy (anonymous):

yay!